### v4

@@ 2 Data and Methods
Psuedo-time assignment, also known as trajectory inference, is the ordering of Single-cell RNA Sequencing experiments into trajectories and steady states. We have developed a novel pesudo-time assignment algorithm that is based on our modeling methodology, and is analogous an expectation-maximization procedure. Initially, the reconstructed network is set to the model that contains all the candidate edges, with some initial logic. For example, if edge signs are available, a rule like ”inhibitors win over activators’ can be used. Next, using the initial network, network trajectories from the states corresponding to the observed Boolean states of the cells are generated: each trajectory is extended until an attractor is reached, and also includes all attractor states. Each observed Boolean cell state is then mapped to its closest state in the trajectories that were obtained. Contiguous sequences of states are assigned to the corresponding trajectories time points, where trajectories of length 1 are set as steady states. Using the assignment of pseudo-time, a new network is reconstructed from the data, and a set of de-noised cell states is obtained, from which a new pseudo-time is generated. Using the new pseudo-time, a network is fitted to the data. The process repeats itself until the value of the solution stops improving. In the EM analogy, the latent variables are the cell assignments to pseudo-time points, and the network edges and logic are the model parameters.
In the next section we demonstrate the effectiveness of our method on two real datasets.
@@ 3 Results
The methodology described above was implemented using the Gurobi python API, which allows for a high level of correspondence between the code and the mathematical description of the model. We set the solver parameter OBBT to the value 3. The straightforward correspondence between the python API and the variables and constraints of the problem will enable other programmers an easy access to our implementation.
@@
To test our method on real experimental data, we obtained the microarray dataset GSE49650 of synchronized yeast cells from the Gene Expression Omnibus. Preprocessing of the .CEL files was done using the rma function in the Bioconductor package affy, using default argument values. Additionally, every trajectory (time-series) of every gene was smoothed using the R functions smooth.spline with smooth parameter 0.5 and then approxfun (in the stats package) with default argument values. The x-coordinate values for the smoothing were the times at which the measurements were taken in minutes, and the y-coordinate values were the array intensities. We used the BASCA method as implemented in the R package Binarize (Hopfensitz et al., 2012), with default arguments, for mapping from continuous to Boolean values. Every trajectory was binarized separately. We used the yeast cell cycle model from Cho et al. (Cho et al., 2019), with the edges as the candidate regulatory connections. Complexes were modeled using the expression of one of their genes. After fitting the model, the percentage of mismatches was 16.64 %.
@@
In Cho et al., edges were associated with activation or repression activity. Compared to the inferred logic tables, all the edges agreed on the sign with Cho et al. This result is detailed in Table 1. The inferred logic provided information about combinatorial regulations, i.e. what is the effect of multiple regulators that regulate that target at the same time. This is illustrated in Figure 1 for two of the targets. Interestingly, genes can be separated into those that are active in early stages and those that are active in later stages (Figure 2). Figure 2b shows an inferred trajectory for one of the time-series, and Figure 2a shows the continuous values that were binarized for two of the genes in the model. It should be noted that not every interaction was inferred. The reasons for this is that either the dataset size is too small, or that the interactions occurring in the dataset do not capture all the interactions in the model. Interestingly, both the synchronous and asynchronous dynamics of the inferred model lead to the same steady state, suggesting a robust design. Perturbing a single gene in the steady state generates an oscillation that settles back to the steady state.
Next, we applied our pseudo-time inference algorithm to the HSC network of Bonzanni et al. (Bonzanni et al., 2013), using GEO dataset GSE75478 which profiles human hematopoeitic stem cells in early differentiation, in order to try to reconstruct the network from single-cell data. The raw data from the file GSE75478_transcriptomics_raw_filtered_I2.csv.gz was preprocessed using the R package scuttle, followed by library-size normalization and log transformation. Binarized values were obtained using the R Binarize package and the kMeans method, with default arguments. As an initial network for the pseudo-time inference algorithm, we used the edge signs published in Bonzanni et al. with the ”inhibitors win” rule for combinatorial interactions. The initial rate of mismatches was 0.158, and after convergence it was reduced to 0.126. All the inferred interactions matched the signs of the edges published in Bonzanni et al., though for the regulation of RUNX1, the activiating effect of its regulators seem to require two active regulators (Table 2). We conclude from both experiment that accurate network reconstruction is possible from varying data types and in the presence of asynchronous dynamics.
[…]
@@
[…]
| GATA1 | SPI1 | repression | repression |
| FLI1 | SPI1 | activation | activation |
| RUNX | SPI1 | activation | activation |
| SPI1 | GATA1 | repression | repression |
| TAL1 | GATA1 | activation | activation |
| GATA1 | TAL1 | activation | activation |
| ERG1 | TAL1 | activation | activation |
| GATA2 | TAL1 | activation | activation |
| GATA1 | FLI1 | repression | repression |
| ERG1 | FLI1 | activation | activation |
| TAL1 | FLI1 | activation | activation |
| RUNX1 | ERG | activation | activation |
| TAL1 | ERG | repression | repression |
| TAL1 | CBFA2T3 | activation | activation |
| SPI1 | RUNX1 | activation | activation |
| TAL1 | RUNX1 | activation | activation |
| GATA2 | RUNX1 | activation | activation |
[…]
@@ 4 Conclusion
In this work we have presented a methodology that chooses an optimal model given a time-series dataset that contains asynchronous updates. Our methodology, implemented in software and publicly available for the community, can greatly enhance researchers’ ability to understand their data in the context of a regulatory network. The experimental datasets that we analyzed exhibit asynchronicity, but nevertheless our algorithm was able to successfully infer the regulatory interactions. We were also able to utilize the inference method in a pseudo-time inference algorithm that makes use of network trajectories. Based on these results, we believe that the algorithm is applicable to a broad range of high-throughput datasets. Several challenges are left for future work: first, the computational efficiency of the method should be further improved, as the problem it addresses is likely to present a variety challenging instances. Another challenge is identifying the best approach for selecting a gene’s candidate regulators out of the total set of network nodes. Finally, the topic of binarization of continuous or discrete data into Boolean values deserves more study. Such work should examine the effects of different binarization techniques on inference of asynchronous dynamics systematically, and their synergy with different experimental data types and preprocessing procedures. Special attention should be given to the effects of different technology, particularly single-cell vs. bulk RNA-Seq.
5 Conflict of Interests
The authors declare that they have no conflicting interests.
6 Availability of data and software code
The code and data are publicly available on GitHub: https://github.com/karleg/MEDSI

### v5

@@ 2.2 Pseudo-time
Psuedo-time assignment, also known as trajectory inference, is the ordering of individual states, produced by scRNA-Seq experiments, into trajectories and steady states. The method described in the previous section can be adopted for this purpose, in a procedure that resembles Expectation-Maximization. Initially, the reconstructed network is set to the model that contains all the candidate edges, with some initial logic. For example, if edge signs are available or estimated from correlations, a rule like "inhibitors win over activators’ can be used. Next, using the initial network, network trajectories from the states corresponding to the observed Boolean states of the cells are generated: each trajectory is extended until an attractor is reached, and also includes all attractor states. Each observed Boolean cell state is then mapped to its closest state in the trajectories that were obtained. Contiguous sequences of states are assigned to the corresponding trajectories time points, where trajectories of length 1 are set as steady states. Using the assignment of pseudo-time, a new network is reconstructed from the data, and a set of de-noised cell states is obtained as part of this reconstruction. Each of these can be used as an initial state for the inferred network in order to generate trajectories, map the data to them, and obtain a new pseudo-time. Using the new pseudo-time, a network is fitted to the data. The process repeats itself until the value of the solution stops improving. In the EM analogy, the latent variables are the pseudo-times assigned to cell states, and the network edges and logic are the model parameters. For ease of reference we will refer to this procedure as TICO (Timeless Inference of Cell Ordering).
@@ 3 Results
3.1 Synchronized Cells
@@
To test our method on real experimental data, we obtained the microarray dataset GSE49650 of synchronized yeast cells from the Gene Expression Omnibus. Preprocessing of the .CEL files was done using the rma function in the Bioconductor package affy, using default argument values. Additionally, every trajectory (time-series) of every gene was smoothed using the R functions smooth.spline with smooth parameter 0.5 and then approxfun (in the stats package) with default argument values [27]. The x-coordinate values for the smoothing were the times at which the measurements were taken in minutes, and the y-coordinate values were the array intensities. We used the BASCA method as implemented in the R package Binarize [21], with default arguments, for mapping from continuous to Boolean values. Every trajectory was binarized separately.
We used the yeast cell cycle model from Cho et al. [28], with the edges as the candidate regulatory connections. Complexes were modeled using the expression of one of their genes. After fitting the model, the percentage of mismatches was 16.3 %.
@@
In Cho et al., edges were associated with activation or repression activity. Compared to the inferred logic tables, all the edges agreed on the sign with Cho et al. This result is detailed in Table 1. The inferred logic provided information about combinatorial regulations, i.e. what is the joint effect of multiple regulators. This is illustrated in Figure 1 for two of the targets. Figure 2 shows an inferred trajectory for one of the time-series, and the original continuous levels for two of the genes. Interestingly, genes can be separated into those that are active at early or late cell cycle stages.
It should be noted that not every candidate interaction was inferred. The reason for this is that either the dataset size is too small for obtaining reliable predictions, or that the interactions occurring in the dataset do not capture all the candidate interactions. Both the synchronous and asynchronous dynamics of the inferred model lead to the same steady state, suggesting a robust design. Perturbing a single gene in the steady state generates an oscillation that settles back to the steady state.
3.2 scRNA-Seq
Next, we applied the pseudo-time inference algorithm (TICO) to the HSC network of Bonzanni et al. [29], using GEO dataset GSE75478 [30], which profiles human hematopoeitic stem cells in early differentiation from the bone marrow of two individuals.
The raw data was preprocessed using the R package scuttle, followed by library-size normalization and log transformation. Binarized values were obtained using the BASCA method [21] from the R package Binarize, with default arguments. As an initial network for the pseudo-time inference algorithm, we used the logic published in Bonzanni et al. For validation, we used the CD38 measurements for individual cells, as CD38 is a marker for differentiation and therefore expected to increase along differentiation trajectories.
The binarized cell states were mapped to their closest counterparts in the network trajectories for the initial and inferred networks, and Pearson correlation and the corresponding correlation test p-values were computed between CD38 measurements in consecutive states. As shown in figure 3, correlation values from the inferred model were higher, with fold changes of 1.7 and 2.6 for individuals 1 and 2, respectively. The corresponding p-values were also lower, and significant for the inferred networks -  for individual 1 and  for individual 2. The  ratio of the p-values obtained for the inferred and the initial networks are displayed in figure 3.
3.3 Simulation
To test TICO using a known ground-truth, we generated 20 simulated datasets from networks the size of Bonzanni et al. Two random true regulators and one random false regulator, and an interaction between the false regulator and true ones such that it has maximal agreement with the data, were added for each gene. In order to generate count data that is similar to scRNA-Seq experiments, we binarized the real data from GEO dataset GSE75478 to obtain active and inactive labels, and fitted each gene with a zero-inflated negative binomial model. Network dynamics were generated using the R package BoolNet [31], with some asynchronous updates, and Boolean values were mapped to counts by generating random values from the fitted models of randomly chosen genes in GSE75478.
For comparison, we used SCORPIUS [32] to infer pseudo-time and the bestfit method from the Boolnet [31] package to reconstruct the networks. Since the pseudo-time generated by SCORPIUS does not assign the cells to individual trajectories, we provided SCORPIUS with the trajectory labels - only the order of cells/states within trajectories needed to be inferred. Additionally, SCORPIUS does not provide directionality in time. Therefore, for SCORPIUS predictions we chose the directionality that agrees best with the true trajectory. These choice provide a significant boost to SCORPIUS predictions - for example, trajectories of length 2 will always be inferred correctly. Nevertheless, the predictions obtained by TICO were significantly better. Figure 4 shows the true positives (y-axis) and false positives (x-axis) for different datasets obtained by the two methods(blue- TICO, red - SCORPIUS+bestfit). The green line is the ratio of TP and FP expected by chance. As can be seen in the figure, the blue dots are significantly removed from the green line, whereas the red ones are spread along its margin. When computing the probability of randomly drawing the observed number of TPs or more, the upper tail of the hypergeometric distribution is less than 0.05 for 15 of reconstructions obtained by TICO, and only for one reconstruction obtained by SCORPIUS+bestfit.
| Source | Target | Cho etl al. | GSE49650 |
| MCM1 | CLN3 | activation | activation |
@@ 4 Conclusion
In this work we presented a novel network-inference methodology that finds an optimal solution with respect to minimizing spurious fits, and can account for asynchronous updates in network dynamics. Our methodology is implemented in software and publicly available for the community. It can greatly enhance researchers’ ability to understand their data in the context of a regulatory network. The experimental datasets that we analyzed exhibit asynchronicity and our algorithm was able to successfully infer the regulatory interactions. Based on this result, we believe that the algorithm is applicable to a broad range of high-throughput datasets. Several objectives are left for future work: First, additional heuristics should be developed and studied, as the problem it addresses is likely to present a variety of challenging instances. Second, the topic of binarization of continuous or discrete data into Boolean values should be further pursued to obtain better understanding of current experimental technologies. Finally, combining different types of biological networks into a single model can provide broader insights into cellular function and should be studied using the method described in this work.
Declarations
4.1 Competing Interests
The authors declare that they have no competing of interests.
4.2 Data Availability
An implementation of the method described in this work can be found at https://github.com/karleg/MEDSI The data is used in this work is available publicly at the Gene Expression Omnibus at https://www.ncbi.nlm.nih.gov/geo/ ==========================================================